OpenStudy (anonymous):

z^(1/3) divided by cube root of z squared

OpenStudy (lυἶცἶ0210):

Is it this? \[\Large \frac{z^{1/3}}{\sqrt[3]{z^2}}\]

OpenStudy (anonymous):

yes

OpenStudy (lυἶცἶ0210):

\(\Large \sqrt[3]{z^2}\) is the same as \(\Large z^{2/3}\), so convert it into that and it's just power fraction rules from there

OpenStudy (anonymous):

Thanks

OpenStudy (lυἶცἶ0210):

\(\Large \frac{x^a}{x^b}=x^{a-b}\) And you're welcome :)